Circular No. 6657

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970508 A. J. Castro-Tirado and J. Gorosabel, Laboratorio de Astrofisica Es-

pacial y Fisica Fundamental, Madrid; C. Wolf and R. Fockenbrock, Max-Planck-Institut für Astronomie (MPIA), Heidelberg; E. Martinez-Gonzalez and N. Benitez, Instituto de Fisica de Cantabria, Santander; J. Greiner, Astrophysikalisches Institut, Potsdam; E. Costa, M. Feroci and L. Piro, Istituto di Astrofisica Spaziale, Frascati; F. Frontera, Università di Ferrara; E. Palazzi, G. Pizzichini, and L. Nicastro, Istituto Tecnologie e Studio Radiazioni Extraterrestri, CNR, Bologna; J. Heise and J. in’t Zand, Space Research Organization of the Netherlands, Utrecht; C. Bartolini, A. Guarnieri, and A. Piccioni, Università di Bologna; N. Masetti, Università di Padova; J. Gallego and J. Zamorano, Universidad Complutense, Madrid; H. Pedersen, Copenhagen Astronomical Observatory; and K. Birkle, MPIA, report that R-band images of the field of GRB 970508 were obtained with the 2.2-m CAHA telescope (+ CAFOS) at Calar Alto on May 9.128, 9.899, 10.872, and 11.868 UT. The variable object noted by Bond (IAUC 6654) yielded respective preliminary magnitudes (± 0.2) R = 20.8, 20.8, 19.6 and 20.4, perhaps indicating the beginning of the decline phase. A 200-s spectrum (range 430–710 nm) obtained on May 10.879 shows a blue continuum with no obvious strong emission lines. A second spectrum (350–800 nm) with 3600-s integration time was obtained on May 11.895, suggesting the absence of strong emission and absorption lines. Images were also obtained in the U band with the 4.2-m William Herschel Telescope at La Palma on May 9.088, 10.056, 11.092, yielding preliminary magnitudes (± 0.3) 21.0, 20.5, and 20.3. The object is therefore very blue and starlike in shape.

COMET 68P/KLEMOLA R-band CCD photometry by C. W. Hergenrother using the 1.2-m re-

fector at the F. L. Whipple Observatory: Mar. 29.41 UT, 21.4; May 3.33, 19.6; 3.34, 19.4. The images are essentially stellar on both nights but within 1′′ of the comet’s expected position.

COMET 103P/HARTLEY 2 R-band CCD photometry by Hergenrother: May 2.47 UT, 19.6; 3.43,

19.5. Co-adding three images (total integration time 900 s) on May 3 shows a coma 5′′ across and a tail 8′′ long in p.a. 255^{o}.

1997 May 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green